# Zyda: A 1.3T Dataset for Open Language Modeling

### v1

@@ 1 Introduction
Open-source models have, to some extent, kept up with the frontier in performance during this period of unprecedented scaling. While these models (Touvron et al.,, 2023; Team et al.,, 2024; Jiang et al.,, 2023) are almost always ’open-weights’, they usually give little to no information about the datasets they are trained on. In part due to this, openly available datasets have tended to lag significantly behind the frontier in both scale and quality.
@@
If open models are to become competitive with the state of the art, they will need extremely large, general, and high-quality open datasets which can be utilized straightforwardly to train such models, allowing practitioners to focus on other aspects of performance such as model scaling and architecture. Moreover, large high-quality datasets may encourage dataset standardization which can allow for fair and meaningful comparisons between different potential architectures and training methods.
@@
In this paper, we take a step towards this vision by releasing Zyda – an open, permissively licensed dataset of 1.3 trillion tokens. Zyda was created by combining major permissively licensed open datasets which are recognized as high-quality within the community. Beyond simple collation, we performed extensive and thorough additional filtering and both intra- and inter-dataset deduplication on these datasets, in addition to any processing they originally underwent.
Specifically, we design, test, and tune a novel filtering pipeline which is significantly more comprehensive than prior works. This pipeline utilizes a wide range of filters from a number of sources (Raffel et al.,, 2020; Soboleva et al.,, 2023; Kudugunta et al.,, 2024; Penedo et al.,, 2023) and also introduces a number of novel ones, all of which were extensively manually tested and tuned to ensure efficacy.
@@
Building atop the deduplication pipeline from (Soboleva et al.,, 2023), we optimize it for highly parallel processing and carefully tune our deduplication hyperparameters with manual inspections to ensure a good false-positive rate. Inter-dataset deduplication is crucial, since simply combining existing datasets does not remove the duplicate data which is found in both datasets independently. We performed cross-dataset deduplication and found large numbers of cross-dataset duplicates, which is not surprising given that almost all existing open-source datasets were ultimately derived from Common Crawl and filtered in similar ways.
@@
We find that models trained on Zyda perform strongly on language modelling tasks, significantly outperforming models trained upon Dolma as well as the Pile (see Fig. 5(a) left panel). Further, we find that StarCoder interferes with language modelling abilities, and when we remove the StarCoder subset, Zyda outperforms all its constituent datasets, including RefinedWeb which is known to be a particularly high quality dataset. While these comparisons are done at equi-token level, Zyda contains double the amount of tokens as RefinedWeb, thus enabling significantly better performance if the entire dataset is trained upon. We believe this improvement is based upon our postprocessing pipeline – especially the heavy deduplication both within and across datasets.
Zyda is openly available on Huggingface at https://huggingface.co/datasets/Zyphra/Zyda and all of our dataset processing code is open-source and is available at https://github.com/Zyphra/Zyda_processing.
@@ 2.2 Filtering
For the second stage, we performed document-level filtering based on a set of syntactic heuristics which can be cheaply computed from the raw text of a document. If the threshold of the filter was exceeded, the whole document was removed from the dataset. These filters broadly fell into three categories: (1) removing syntactically broken or otherwise gibberish documents, (2) removing semantically meaningless documents, and (3) removing meaningful but objectionable content. Examples of the first kind of filter include methods like filtering based on the proportion of alphanumeric characters in a document, which at high proportions uniformly corresponds to documents comprised of entirely gibberish strings generated either through broken preprocessing or through unknown processes on the internet. Semantically meaningless documents include documents full of seemingly random numbers, cryptographic strings, and lists of seemingly unrelated URLs. Objectionable content included primarily pornographic and offensive content, which we removed with specialized word lists.
Our general philosophy in filtering was to not filter excessively and keep false positives (i.e. good documents removed) relatively low. We manually tested and tuned each of the filters presented here on the Pile dataset. We tuned the filter thresholds so that we obtained a false-positive threshold of about 20% – that is, 20% of the filtered documents were seemingly unobjectionable, while 80% were obviously harmful and it was correct to remove them. This provides a reasonable trade-off between excessive filtering while still removing the majority of content the filter was aimed at reducing. For every filter, we spent significant time manually tuning the threshold and other parameters, then looking at the filtered outputs and attempting to identify true vs false positives. We primarily performed this tuning on the Pile dataset and performed only sanity checking of the filter outputs on the other datasets. However, we believe that many of the categories of low-quality text likely have similar distributions between datasets, since so many of them are derived from the same source. A full list and description of all filters can be found in Appendix B.1.
All datasets were filtered using the same rules except for StarCoder, which we exempted since it consists entirely of code which has a significantly different distribution than the primarily text data in the other datasets. Additionally, the StarCoder authors performed a thorough code-specific filtering of the dataset before release. Table 1 shows the proportion and number of rows removed from each dataset by our filtering process. We observe that the primary source of data removed by our filters is the Pile and arXiv.
@@ 2.3 Deduplication
Since the LSH minhash algorithm only performs approximate deduplication, we also sampled 4.8 million duplicate pairs to estimate the actual false-positive rate based on Jaccard and edit similarities. We define the edit similarity between two documents as their edit distance divided by maximum length of two documents. We found good agreements between theoretical and estimated false-positive rate based on Jaccard similarities, while the estimated false-positive rate based on edit similarity for the 40% threshold was even lower than the theoretical estimate at $3.1\%$. Figures 2(a) and 2(b) show the distribution of edit and Jaccard similarities for the 40% threshold version of our dataset: As expected, the vast majority of identified duplicates are above the threshold (marked as red dash line), and for the edit similarity metric, the distributions is skewed toward higher values (which is expected, since it corresponds to a lower false-positive rate).
To compare the 40% and the 80% versions of Zyda, we trained 1.4B transformers on 50B tokens sampled from each version of Zyda. We found that 40% Zyda performed slightly better (Fig. 5(b)). Due to this, we chose to release the 40% version as our primary Zyda dataset.
@@ 3 Performance
Next we perform a preliminary scaling analysis of our dataset against the Pile and Pythia on a range of scales from 160M to 1.4B, all trained on the full Pile and 300B tokens of Zyda. We show that the advantage of Zyda appears to increase with scale. We suggest that this is because the increased dataset quality is not ‘visible’ to small models as they are capacity limited to only extract the larger modes of variance in the data, whereas larger models can additionally ‘see’ finer modes of variance where our increased dataset quality provides an advantage. A small model may simply be unable to fit important aspects of language regardless of the dataset used to train it.
Finally, we performed a number of ablations of our dataset and comparisons to other datasets at the 50B tokens, 1.4B model scale. We plot the aggregate performance on a range of evaluations of final checkpoints on Figure 5(a), and performance as a function training steps on Figure 3.
@@ 4 Related Work
We built our deduplication pipeline upon SlimPajama’s existing open-source libraries, although we ended up significantly modifying them and performed extensive manual tuning and optimization. Our initial filters were inspired by the filter list accompanying the RedPajama dataset. However, after preliminary testing, we removed some of their filters which we judged were ineffective and integrated new ones, either from other papers such as C4, Gopher, and MADLAD-400 (Kudugunta et al.,, 2024), or of our own invention. A full list and description of the filters we used, as well as qualitative impressions gained during testing them, can be found in Appendix B.1.
@@ 5 Discussion
Appendix A Ablation performance by evals
@@
Appendix B Additional details for dataset processing
B.1 Additional Filtering Details
@@
B.2 Number of documents removed by each filter per dataset
@@ B.3 Duplicates sources
B.4 Estimating empirical false positive rate
@@
Figure 8: Statistics of the sample of 4.8 duplicate documents.

### v2

@@ 1 Introduction
Open-source models have, to some extent, kept up with the frontier in performance during this period of unprecedented scaling. While these models (Touvron et al.,, 2023; Team et al.,, 2024; Jiang et al.,, 2023) are almost always ’open-weights’, they usually give little to no information about the datasets they are trained on. In part due to this, openly available datasets have tended to lag significantly behind the frontier in both scale and quality. If open models are to become competitive with the state of the art, they will need extremely large, general, and high-quality open datasets which can be utilized straightforwardly to train such models, allowing practitioners to focus on other aspects of performance such as model scaling and architecture. Moreover, large high-quality datasets may encourage dataset standardization which can allow for fair and meaningful comparisons between different potential architectures and training methods.
@@
In this paper, we take a step towards this vision by releasing Zyda – an open, permissively licensed dataset of 1.3 trillion tokens. Zyda was created by combining major permissively licensed open datasets which are recognized as high-quality within the community. Beyond simple collation, we performed extensive and thorough additional filtering and both intra- and inter-dataset deduplication on these datasets, in addition to any processing they originally underwent. Specifically, we design, test, and tune a novel filtering pipeline which is significantly more comprehensive than prior works. This pipeline utilizes a wide range of filters from a number of sources (Raffel et al.,, 2020; Soboleva et al.,, 2023; Kudugunta et al.,, 2024; Penedo et al.,, 2023) and also introduces a number of novel ones, all of which were extensively manually tested and tuned to ensure efficacy.
@@
Building atop the deduplication pipeline from (Soboleva et al.,, 2023), we optimize it for highly parallel processing and carefully tune our deduplication hyperparameters with manual inspections to ensure a good false-positive rate. Inter-dataset deduplication is crucial, since simply combining existing datasets does not remove the duplicate data which is found in both datasets independently. We performed cross-dataset deduplication and found large numbers of cross-dataset duplicates, which is not surprising given that almost all existing open-source datasets were ultimately derived from Common Crawl [^1] https://commoncrawl.org/ and filtered in similar ways.
@@
We find that models trained on Zyda perform strongly on language modelling tasks, significantly outperforming models trained upon Dolma as well as the Pile (see Fig. 5(a) left panel). Further, we find that StarCoder interferes with language modelling abilities, and when we remove the StarCoder subset, Zyda outperforms all its constituent datasets, including RefinedWeb which is known to be a particularly high quality dataset. While these comparisons are done at equi-token level, Zyda contains double the amount of tokens as RefinedWeb, thus enabling significantly better performance if the entire dataset is trained upon. We believe this improvement is based upon our postprocessing pipeline – especially the heavy deduplication both within and across datasets. Zyda is openly available on Huggingface at https://huggingface.co/datasets/Zyphra/Zyda and all of our dataset processing code is open-source and is available at https://github.com/Zyphra/Zyda_processing.
@@ 2.2 Filtering
Our general philosophy in filtering was to not filter excessively and keep false positives (i.e. good documents removed) relatively low. We manually tested and tuned each of the filters presented here on the Pile dataset. We tuned the filter thresholds so that we obtained a false-positive threshold of about 20% – that is, 20% of the filtered documents were seemingly unobjectionable, while 80% were obviously harmful and it was correct to remove them. This provides a reasonable trade-off between excessive filtering while still removing the majority of content the filter was aimed at reducing. For every filter, we spent significant time manually tuning the threshold and other parameters, then looking at the filtered outputs and attempting to identify true vs false positives. We primarily performed this tuning on the Pile dataset and performed only sanity checking of the filter outputs on the other datasets. However, we believe that many of the categories of low-quality text likely have similar distributions between datasets, since so many of them are derived from the same source. A full list and description of all filters can be found in Appendix E.1. All datasets were filtered using the same rules except for StarCoder, which we exempted since it consists entirely of code which has a significantly different distribution than the primarily text data in the other datasets. Additionally, the StarCoder authors performed a thorough code-specific filtering of the dataset before release. Table 1 shows the proportion and number of rows removed from each dataset by our filtering process. We observe that the primary source of data removed by our filters is the Pile and arXiv.
@@ 2.3 Deduplication
Since the LSH minhash algorithm only performs approximate deduplication, we also sampled 4.8 million duplicate pairs to estimate the actual false-positive rate based on Jaccard and edit similarities. We define the edit similarity between two documents as their edit distance divided by maximum length of two documents. We found good agreements between theoretical and estimated false-positive rate based on Jaccard similarities, while the estimated false-positive rate based on edit similarity for the 40% threshold was even lower than the theoretical estimate at $3.1\%$. Figures 2(a) and 2(b) show the distribution of edit and Jaccard similarities for the 40% threshold version of our dataset: As expected, the vast majority of identified duplicates are above the threshold (marked as red dash line), and for the edit similarity metric, the distributions is skewed toward higher values (which is expected, since it corresponds to a lower false-positive rate). To compare the 40% and the 80% versions of Zyda, we trained 1.4B transformers on 50B tokens sampled from each version of Zyda. We found that 40% Zyda performed slightly better (Fig. 5(b)). Due to this, we chose to release the 40% version as our primary Zyda dataset.
@@ 3 Performance
Next we perform a preliminary scaling analysis of our dataset against the Pile and Pythia on a range of scales from 160M to 1.4B, all trained on the full Pile and 300B tokens of Zyda. We show that the advantage of Zyda appears to increase with scale. We suggest that this is because the increased dataset quality is not ‘visible’ to small models as they are capacity limited to only extract the larger modes of variance in the data, whereas larger models can additionally ‘see’ finer modes of variance where our increased dataset quality provides an advantage. A small model may simply be unable to fit important aspects of language regardless of the dataset used to train it. Finally, we performed a number of ablations of our dataset and comparisons to other datasets at the 50B tokens, 1.4B model scale. We plot the aggregate performance on a range of evaluations of final checkpoints on Figure 5(a), and performance as a function training steps on Figure 3.
@@ 4 Related Work
We built our deduplication pipeline upon SlimPajama’s existing open-source libraries, although we ended up significantly modifying them and performed extensive manual tuning and optimization. Our initial filters were inspired by the filter list accompanying the RedPajama dataset. However, after preliminary testing, we removed some of their filters which we judged were ineffective and integrated new ones, either from other papers such as C4, Gopher, and MADLAD-400 (Kudugunta et al.,, 2024), or of our own invention. A full list and description of the filters we used, as well as qualitative impressions gained during testing them, can be found in Appendix E.1.
@@ 5 Discussion
Appendix A Limitations
While we believe Zyda is a highly effective dataset for language modelling, and this is supported by our ablation studies, and our empirical successes with the Zamba model (Glorioso et al.,, 2024). However, except our ablations comparisons against the Pythia suite, all our other ablations were only trained on 50 billion tokens which may not be enough to predict dataset quality over the full course of training over one or more epochs on the dataset. We limited our ablations to 50B tokens on a 1.4B model primarily out of compute and cost considerations. Additionally, while we performed a small scaling suite for the Pythia models, we did not perform this for other ablations, focusing on the 1.4B scale. It is possible that the scaling properties of different datasets differ so tht some datasets may perform better with larger or smaller models, and we cannot discount that this could change the rank ordering of datasets by performance.
Appendix B Ablation experimental details
For our 50B ablations, we used a 1.4B standard transformer model. The architecture and trainng hyperparameters are presented in Table 3. We used the Swiglu activation function on the MLP layers and trained with RoPE position embeddings.
Hyperparameter Value
Number of Layers 24
Hidden Dimension 248
Number of Attention Heads 16
Batch Size 512
Max Learning Rage 2.0e-4
LR Decay Schedule Cosine
Minimum LR 2.0e-5
Weight Decay 0.1
Adam Beta2 0.95
LR Warmup 0.01
Gradient Clipping 1.0
Training Precision BF16
Activation Function Swiglu
Table 3: Model and training hyperparameters for the 1.4B transformer models trained in our ablation experiments.
For the Pythia comparisons, we trained on 300B tokens (the same size as full Pile), using the model and training hyperparameters reported in Biderman et al., (2023).
All models were trained on Nvidia H100 DGX systems. Most ablations were run on 1-2 DGX nodes. We used the MegatronLM (Shoeybi et al.,, 2019) framework for training the models used in these ablation studies.
Appendix C Societal impacts
We release Zyda with the aim of providing a dataset to make language model training at the trillion-token scale more accessible and effective. Empowering disparate and less-resourced groups to create such language models may have a variety of societal impacts. While enabling more actors to train effective language models may lead to greater decentralization of power and more equal access to the technology, it also opens the door to potential misuse. While we have endeavoured with our filtering to remove harmful and personally identifying content from Zyda, we cannot guarantee that Zyda cannot be used to train language models which could cause harm.
Appendix D Ablation performance by evals
@@
Appendix E Additional details for dataset processing
E.1 Additional Filtering Details
@@
E.2 Number of documents removed by each filter per dataset
@@ E.3 Duplicates sources
E.4 Estimating empirical false positive rate
@@
Figure 8: Statistics of the sample of 4.8 million duplicate documents.